OpenStudy (anonymous):

finding the anti derivatives by substitution. 2/(1-x)^2dx

OpenStudy (anonymous):



OpenStudy (anonymous):

oh parenthesis around the (1-x)^2

OpenStudy (anonymous):

I have u=1-x du=d(1-x) du=-1 dx

zepdrix (zepdrix):

good, from there, multiply both sides by -1 in your last step

zepdrix (zepdrix):

\[\Large dx=-du \qquad\text{right?}\]